# DeFi Arbitrage in Hedged Liquidity Tokens

### v2

@@ 1 Introduction
Similar studies have been undertaken previously that highlight different aspects which we will consider. For example [19, 14] recognize that the liquidity tokens payoff behave like that of a call option. However, as far as we are aware, our approach of expressing this optionality characteristic via the volatility of the price process and defining the implied volatility so as to match the price is wholly novel in the literature. For instance, [11] presents an approximating static hedge for impermanent loss in the CPMM using variance and gamma swaps. This approach is extended in [17] to consider the concentrated liquidity framework of Uniswap v3. In contrast, [19] introduces the loss-versus-rebalancing which considers the cost of dynamic replication of the underlying pool of assets for generic AMMs. Similarly, [7] introduces a decomposition of this dynamic replication of the underlying pool of assets which allows the introduction of a so-called “predictable loss”. Finally, [9] considers an empirical pricing of a liquidity token in an AMM using historically calibrated parameters. Herein, rather than concentrating on the underlying holdings of the CPMM, we will consider the stream of fees as the primary driver of the value of the liquidity token. We note that [16] considers a different approach to estimate the expected fees collected by a LP in a CPMM.
@@
The rest of this paper is organized as follows. In Section 2, we provide a brief introduction to the mathematics of AMMs and, applying this construction to Uniswap data, we explore the pricing of a CPMM liquidity token in practice. Notably, within Example 2.3, we find that the studied data readily admits arbitrage opportunities. With this motivation, in Section 3, we provide the main mathematical theory for risk-neutral pricing and hedging the liquidity token of a CPMM. In Section 4, we provide mathematical discussions surrounding the market implied volatility and a scheme to calibrate a fair valuation of the liquidity token that does not readily admit arbitrage opportunities. In doing so, we revisit Example 2.3 and apply the risk-neutral pricing theory to the data so as to calibrate an arbitrage-free pricing of the Uniswap liquidity token. Finally, in Section 5, we provide novel AMM designs which would eliminate these arbitrage opportunities. In proposing these new constructions, we emphasize potential drawbacks which could occur if implemented in practice.
@@ 2.2 Hedging a CPMM
Consider the CPMM $u(x,y)=xy$. By construction, its pricing oracle $P(x,y)=y/x$ is given by the ratio of the pool’s asset holdings. Due to this construction, LPs mint new liquidity tokens (i.e., deposit) by providing assets at the same ratio as the pool currently holds. Uniswap v2 [2] defines the number of outstanding liquidity tokens to be $L=\sqrt{xy}$ where $(x,y)$ is are the shares of assets held by the pool. Notably, for the CPMM, there is a one-to-one relation between the asset holdings of the pool $(x,y)$ and, jointly, the price $P$ and the number of liquidity tokens $L$. Already we have provided how $P,L$ are constructed from the asset holdings; conversely, given the price and liquidity tokens: $x=L/\sqrt{P}$ and $y=L\sqrt{P}$. In fact, these relations make clear that the value held in the pool is given by $Px+y=2L\sqrt{P}$; because of the linear relation between the assets and the liquidity tokens, throughout the remainder of this work we consider the value of a single liquidity token.
Given this value of the liquidity token, we can eliminate entirely the riskiness of the position by trading the underlying token appropriately (see, e.g., the discussion in and preceding [15, Chapter 2.4]). Specifically, we can hedge this position by holding $\Delta$ units of the underlying where $\Delta$ is the sensitivity of the position to the token price, i.e., $\Delta=\frac{d}{dP}2\sqrt{P}=1/\sqrt{P}$. In theory, by purchasing and rebalancing this position, we expect to perfectly hedge the position. In the following numerical example, we take data on a Uniswap pool in order to look for any arbitrage opportunities, i.e., for trends in the hedged position. Herein, we consider the hedging and rebalancing strategy under negligible fees as is taken in the Black-Scholes framework.
@@
Consider a USDC/WETH Uniswap v3 pool on the Polygon blockchain ($\Delta t=2$ seconds) with $\gamma=5bps$ fees between June 1 and June 30, 2023 with the investment made over the entire price line to mimic the CPMM. [^2] This data is taken from smart contract 0x45dda9cb7c25131df268515131f647d726f50608. This pool was chosen as it is a representative Uniswap v3 pool with high liquidity. For this example, we will assume that the risk-free rate $r=5\%$ (annualized) throughout the period of study. In Figure 1(a), we show the discounted values of both a single liquidity token (i.e., $2\sqrt{P}$ plus the collected fees) and of the delta hedged position (with rebalancing of the delta hedge every block). As expected, the volatility of the hedged position is significantly lower than that of the unhedged position. However, as made clear in Figure 1(b), the delta hedged position has a distinct positive trend line indicating the presence of an arbitrage opportunity.
@@
Figure 1: Example 2.3: Comparison of the discounted values of a liquidity token and its delta hedged position in June 2023.
Before continuing, we wish to explore the possibility that the selected risk-free rate $r=5\%$ is the source of this arbitrage. (All parameters besides the risk-free rate $r$ are specified by the Polygon blockchain or Uniswap contract.) As evidenced in Figure 2, any non-negative choice of risk-free rate $r\geq 0$ leads to a comparable arbitrage as found under $r=5\%$. Notably, the arbitrage appears to be monotonic with $r$ with the smallest (but still significant) arbitrage observed with $r=0\%$. [^3] Though not displayed in Figure 2, numerical tests with negative interest rates show that the monotonicity w.r.t. risk-free rate does not continue holding. As a consequence, negative interest rates are also unable to recover an arbitrage-free environment for the liquidity token either. Therefore, the presence of these arbitrage opportunities implies a mispricing in the liquidity token by Uniswap. Within the rest of this work, we will investigate the possible mispricing of the liquidity token so as to create a full theory for pricing and hedging a CPMM.
(a) Varying $r\in[0\%~{},~{}10\%]$.
(b) Varying $r\in[0\%~{},~{}50\%]$
Figure 2: Example 2.3: Comparison of the (discounted) delta hedged liquidity token under varying interest rate environments.
@@ 3.1 Market Model
For the remainder of this work, we will assume that the price process $(P_{t})$ follows the risk-neutral geometric Brownian motion
@@
for risk-free rate $r\geq 0$, volatility $\sigma>0$, Brownian motion $W$ on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t}^{W})_{t\geq 0},\mathbb{P})$, with the filtration $(\mathcal{F}_{t}^{W})_{t\geq 0}$ generated by the Brownian motion, and initial value $P_{0}$. We also assume that the price process $(P_{t})$ is observed, and can be traded, at a secondary market in continuous time, but can only be traded at $\Delta t$ intervals on the blockchain. When trading is possible in both markets simultaneously, the trading happens at the price $P$ and there is no arbitrage between the two markets.
@@
Following Assumption 3.1, the measure $\mathbb{P}$ is a risk-neutral measure. In fact, due to the completeness of the constructed market, $\mathbb{P}$ is the unique risk-neutral measure. [^4] We refer the interested reader to [13, 8] for the discussion on the Fundamental Theorems of Asset Pricing and to [15, Chapters 1 and 2] that show the existence of a hedge in a complete market driven by a Brownian Motion as well as pricing by replication in such a market. Throughout the remainder of this work, we will exploit this fact in order to consider the pricing and hedging of a liquidity token in a CPMM.
@@
Implicitly, when constructing the fees, we are assuming there is only one transaction in each block and it perfectly aligns the price of the CPMM to $P$. When studying the valuation of the CPMM liquidity token in Section 3.2 below, this assumption guarantees that we find a lower bound on the risk-neutral value as other (uninformed) trades may also occur; such trades increase the fees collected by the LPs without altering the fundamental price process.
@@ 3.2 Risk-Neutral Valuation
$\hat{\gamma}\geq\hat{\gamma}^{*}:=2\left[-1+\frac{\Phi\left(\frac{(r+\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma}\right)-e^{-r\Delta t}\Phi\left(\frac{(r-\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma}\right)}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\right]^{-1},$
@@
It remains to verify that this ansatz strategy is optimal. First, assume $\hat{\gamma}\geq\hat{\gamma}^{*}$ so that $V_{0}(P_{0})=\tilde{V}_{0}(P_{0})$. By construction, we note that $\tilde{V}_{0}(P_{0})$ coincides with its continuation value, that is, $\tilde{V}_{0}(P_{0})$ $=e^{-r\Delta t}\mathbb{E}[\tilde{V}_{0}(P_{0}e^{(r-\frac{\sigma^{2}}{2})\Delta t+\sigma Z\sqrt{\Delta t}})+\hat{\gamma}F(P_{0},P_{0}e^{(r-\frac{\sigma^{2}}{2})\Delta t+\sigma Z\sqrt{\Delta t}})]$ for every $P_{0}>0$. Indeed,
@@ 3.3 Greeks
$\frac{\partial}{\partial P}V_{0}(P)=\frac{\hat{\gamma}}{\hat{\gamma}^{*}\sqrt{P}}=\frac{V_{0}(P)}{2P}$
@@
$\frac{\partial^{2}}{\partial P^{2}}V_{0}(P)=-\frac{\hat{\gamma}}{2\hat{\gamma}^{*}P^{3/2}}=-\frac{V_{0}(P)}{4P^{2}}$
@@
$\frac{\partial}{\partial\sigma}V_{0}(P)=\frac{\hat{\gamma}\sqrt{P}e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\left[\sqrt{\frac{\Delta t}{2\pi}}e^{-\frac{r^{2}\Delta t}{2\sigma^{2}}}-\frac{\sigma\Delta t}{4}\frac{\Phi(\frac{(r+\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma})-e^{-r\Delta t}\Phi(\frac{(r-\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma})}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\right]$
for any price $P>0$. In contrast to the delta and gamma of this position, the vega does not have a constant sign. In particular, $\frac{\partial}{\partial\sigma}V_{0}(P)>0$ for $\sigma>0$ sufficiently small and $\frac{\partial}{\partial\sigma}V_{0}(P)<0$ for $\sigma>0$ sufficiently large. Therefore, in contrast to the typical derivatives contracts, e.g., a European call option, the liquidity token has a complex dependency on volatility rather than simply being a long position in volatility. We demonstrate an example of this complex dependency of vega on the volatility in Figure 3(b). Intuitively, when $\sigma>0$ is small, a tiny increase in volatility will lead to more trading and, therefore, fees collected by the LPs, i.e., a positive vega. On the other hand, when $\sigma>0$ is already very large, a further increase in volatility will, also, increase the probability of the price collapsing; this leads to a drop in the price of the token $P$ which results in a negative vega.
@@
Figure 3: Valuation and vega with $P_{0}=1$, $\gamma=5bps$ (i.e., $\hat{\gamma}=5.0025bps$), $r=0.05\%$ (annualized) and $\Delta t=2$ seconds.

### v3

@@ 1 Introduction
Similar studies have been undertaken previously that highlight different aspects which we will consider. For example [21, 15] recognize that the liquidity tokens payoff behave like that of a call option, with the former also applying the idea of hedging and rebalancing for AMMs in continuous time. The approach of those works was extended to discrete time in [20] and further generalized in [22]. However, as far as we are aware, our approach of expressing this optionality characteristic via the volatility of the price process and defining the implied volatility so as to match the price is wholly novel in the literature. For instance, [13] presents an approximating static hedge for impermanent loss in the CPMM using variance and gamma swaps. This approach is extended in [18] to consider the concentrated liquidity framework of Uniswap v3. In contrast, [21] introduces the loss-versus-rebalancing which considers the cost of dynamic replication of the underlying pool of assets for generic AMMs; [7] uses the loss-versus-rebalancing to quote a risk-neutral expected fees and deduces implications of that construction. Similarly, [9] introduces a decomposition of this dynamic replication of the underlying pool of assets which allows the introduction of a so-called “predictable loss”. Finally, [11] considers an empirical pricing of a liquidity token in an AMM using historically calibrated parameters. Herein, rather than concentrating on the underlying holdings of the CPMM, we will consider the stream of fees as the primary driver of the value of the liquidity token. We note that [17] considers a different approach to estimate the expected fees collected by a LP in a CPMM.
@@
The rest of this paper is organized as follows. In Section 2, we provide a brief introduction to the mathematics of AMMs and, applying this construction to Uniswap data, we explore the pricing of a CPMM liquidity token in practice. Notably, within Example 2.3, we find that the studied data readily admits arbitrage opportunities. With this motivation, in Section 3, we provide the main mathematical theory for risk-neutral pricing and hedging the liquidity token of a CPMM. In Section 4, we provide mathematical discussions surrounding the market implied volatility and a way to calibrate a fair valuation of the liquidity token that does not readily admit arbitrage opportunities. In doing so, we revisit Example 2.3 and apply the risk-neutral pricing theory to the data so as to calibrate an arbitrage-free pricing of the Uniswap liquidity token. Finally, in Section 5, we provide novel AMM designs which would eliminate these arbitrage opportunities. In proposing these new constructions, we emphasize potential drawbacks which could occur if implemented in practice.
@@ 2.2 Hedging a CPMM
Consider the CPMM $u(x,y)=xy$. By construction, its pricing oracle $P(x,y)=y/x$ is given by the ratio of the pool’s asset holdings. Due to this construction, LPs mint new liquidity tokens (i.e., deposit) by providing assets at the same ratio as the pool currently holds. Uniswap v2 [2] defines the number of outstanding liquidity tokens to be $L=\sqrt{xy}$ where $(x,y)$ is are the shares of assets held by the pool. Notably, for the CPMM, there is a one-to-one relation between the asset holdings of the pool $(x,y)$ and, jointly, the price $P$ and the number of liquidity tokens $L$. Already we have provided how $P,L$ are constructed from the asset holdings; conversely, given the price and liquidity tokens: $x=L/\sqrt{P}$ and $y=L\sqrt{P}$. In fact, these relations make clear that the value held in the pool is given by $Px+y=2L\sqrt{P}$.
Remark 1.
Because of the linear relation between the value of the pool and the value of a liquidity position (i.e., scaled by $L$), throughout the remainder of this work we consider the value of a single liquidity token $L=1$ except where otherwise explicitly stated.
Given this value of the liquidity token, we can eliminate entirely the riskiness of the position by trading the underlying token appropriately (see, e.g., the discussion in and preceding [16, Chapter 2.4]). Specifically, we can hedge this position by holding $\Delta$ units of the underlying where $\Delta$ is the sensitivity of the position to the token price, i.e., $\Delta=\frac{d}{dP}2\sqrt{P}=1/\sqrt{P}$. In theory, by purchasing and rebalancing this position, we expect to perfectly hedge the position; however, this hedge may fail during periods of extreme volatility as delta hedging only provides protections along the linearly approximated portfolio. In the following numerical example, we take data on a Uniswap pool in order to look for any arbitrage opportunities, i.e., for trends in the hedged position. Surprisingly, given the theory on delta hedging, we find opportunities for risk-free profits with frequent hedges. To preview future results, in the next section we provide a theory of risk-neutral pricing of a liquidity token and deduce the Greeks of this liquidity position; these results are summarized in Table 1. Herein, we consider the hedging and rebalancing strategy under negligible fees as is taken in the Black-Scholes framework.
@@
Consider a USDC/WETH Uniswap v3 pool on the Polygon blockchain ($\Delta t=2$ seconds) with $\gamma=5bps$ fees between market creation (December 20, 2021) and June 21, 2025 with the investment made over the entire price line to mimic the CPMM. [^2] This data is taken from smart contract 0x45dda9cb7c25131df268515131f647d726f50608. This pool was chosen as it is a representative Uniswap v3 pool with high liquidity. For this example, we will assume that the risk-free rate $r=0\%$ throughout the period of study for simplicity and to match the rate earned on USDC. In Figure 1(a), we show the discounted values of both a 1 USDC liquidity position (i.e., $\sqrt{P}/\sqrt{P_{0}}$ plus the collected fees) and of the delta hedged positions (with rebalancing of the delta hedge at different frequencies ranging from every block to every day). As expected, the volatility of the hedged position is significantly lower than that of the unhedged position. However, as made clear in Figure 1(b), when delta is hedged every block, then the hedged position has a distinct positive trend line indicating the presence of an arbitrage opportunity except on distinct dates in which the hedging fails due to large intra-block price swings. We hypothesize that a better hedging strategy (e.g., Gamma hedging) is needed to reduce these jumps. Notably, from February 2022 to March 2024, there is a nearly unbroken stretch of profitable arbitrage opportunities by implementing this delta hedging strategy. As we reduce the hedging frequency, we find a positive trend exists for rehedging every minute, though the profitability disappears for hedging every hour or day.
@@
Figure 1: Example 2.3: Comparison of the discounted values of a 1 USDC investment in liquidity tokens and its delta hedged position from market creation (December 20, 2021) until June 21, 2025.
Before continuing, we wish to expand on the sensitivity of this arbitrage to the chosen risk-free rate $r=0\%$. (All parameters besides the risk-free rate $r$ are specified by the Polygon blockchain or Uniswap contract.) Notably, as the risk-free rate increases, the profitability of this delta hedging strategy decreases due to the time-value of money that is being forfeited by holding USDC. This leads us to conjecture that the profitability would return if a tokenized money market account was used in place of USDC.
@@ 3 Constant Product Market Making
Valuation $V_{0}(P)$ $\frac{2\hat{\gamma}\sqrt{P}}{\hat{\gamma}^{*}}$
Delta $\frac{\partial}{\partial P}V_{0}(P)$ $\frac{\hat{\gamma}}{\hat{\gamma}^{*}\sqrt{P}}$
Gamma $\frac{\partial^{2}}{\partial P^{2}}V_{0}(P)$ $\frac{\hat{\gamma}}{2\hat{\gamma}^{*}P^{3/2}}$
Vega $\frac{\partial}{\partial\sigma}V_{0}(P)$ $\frac{\hat{\gamma}\sqrt{P}e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\left[\sqrt{\frac{\Delta t}{2\pi}}e^{-\frac{r^{2}\Delta t}{2\sigma^{2}}}-\frac{\sigma\Delta t}{4}\frac{\Phi(\frac{(r+\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma})-e^{-r\Delta t}\Phi(\frac{(r-\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma})}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\right]$
Table 1: Summary table of valuation (Theorem 3.2) and Greeks (Section 3.3) of a single liquidity token in a CPMM when $\hat{\gamma}\geq\hat{\gamma}^{*}$ as defined in (3.4).
@@ 3.1 Market Model
For the remainder of this work, following [8], we will assume that the price process $(P_{t})$ follows the risk-neutral geometric Brownian motion
@@
for risk-free rate $r\geq 0$, volatility $\sigma>0$, Brownian motion $W$ on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t}^{W})_{t\geq 0},\mathbb{P})$, with the filtration $(\mathcal{F}_{t}^{W})_{t\geq 0}$ generated by the Brownian motion, and initial value $P_{0}$.
@@
As highlighted within Assumption 3.1, the price process considered herein is consistent with that in [8]. Specifically, that work considers a terminal, log-normally distributed price and calculate the resulting equilibrium from a system of optimizing agents (traders and LPs). While [8] considers a static setting, we extend to the dynamic setting following the same general direction: we assume traders actualize this log-normal distribution in the AMM.
@@
Following Assumption 3.1, the measure $\mathbb{P}$ is a risk-neutral measure. In fact, due to the completeness of the constructed market, $\mathbb{P}$ is the unique risk-neutral measure. [^3] We refer the interested reader to [14, 10] for the discussion on the Fundamental Theorems of Asset Pricing and to [16, Chapters 1 and 2] that show the existence of a hedge in a complete market driven by a Brownian Motion as well as pricing by replication in such a market. Throughout the remainder of this work, we will exploit this fact in order to consider the pricing and hedging of a liquidity token in a CPMM.
Remark 4.
Implicitly, when constructing the fees, we are assuming there is only one transaction in each block to align the price of the CPMM to $P$. When studying the valuation of the CPMM liquidity token in Section 3.2 below, this assumption guarantees that we find a lower bound on the risk-neutral value as other (uninformed) trades may also occur; such trades increase the fees collected by the LPs without altering the fundamental price process.
@@ 3.2 Risk-Neutral Valuation
$\hat{\gamma}\geq\hat{\gamma}^{*}:=2\left[-1+\frac{\Phi\left(\frac{(r+\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma}\right)-e^{-r\Delta t}\Phi\left(\frac{(r-\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma}\right)}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\right]^{-1},$ (3.4)
@@
It remains to verify that this ansatz strategy is optimal. First, assume $\hat{\gamma}\geq\hat{\gamma}^{*}$ so that $V_{0}(P_{0})=\tilde{V}_{0}(P_{0})$. By construction, we note that $\tilde{V}_{0}(P_{0})$ coincides with its continuation value, that is, $\tilde{V}_{0}(P_{0})$ $e^{-r\Delta t}\mathbb{E}[\tilde{V}_{0}(P_{0}e^{(r-\frac{\sigma^{2}}{2})\Delta t+\sigma Z\sqrt{\Delta t}})+\hat{\gamma}F(P_{0},P_{0}e^{(r-\frac{\sigma^{2}}{2})\Delta t+\sigma Z\sqrt{\Delta t}})]$ for every $P_{0}>0$. Indeed,
@@ 3.3 Greeks
$\frac{\partial}{\partial P}V_{0}(P)=\frac{\hat{\gamma}}{\hat{\gamma}^{*}\sqrt{P}}=\frac{V_{0}(P)}{2P}$ (3.6)
@@
$\frac{\partial^{2}}{\partial P^{2}}V_{0}(P)=-\frac{\hat{\gamma}}{2\hat{\gamma}^{*}P^{3/2}}=-\frac{V_{0}(P)}{4P^{2}}$ (3.7)
@@
$\!\!\!\!\!\!\!\!\!\frac{\partial}{\partial\sigma}V_{0}(P)=\frac{\hat{\gamma}\sqrt{P}e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\left[\sqrt{\frac{\Delta t}{2\pi}}e^{-\frac{r^{2}\Delta t}{2\sigma^{2}}}-\frac{\sigma\Delta t}{4}\frac{\Phi(\frac{(r+\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma})-e^{-r\Delta t}\Phi(\frac{(r-\frac{\sigma^{2}}{2})\sqrt{\Delta t}}{\sigma})}{1-e^{-\frac{1}{2}(r+\frac{\sigma^{2}}{4})\Delta t}}\right]$ (3.8)
for any price $P>0$. In contrast to the delta and gamma of this position, the vega does not have a constant sign. In particular, $\frac{\partial}{\partial\sigma}V_{0}(P)>0$ for $\sigma>0$ sufficiently small and $\frac{\partial}{\partial\sigma}V_{0}(P)<0$ for $\sigma>0$ sufficiently large. Therefore, in contrast to the typical derivatives contracts, e.g., a European call option, the liquidity token has a complex dependency on volatility rather than simply being a long position in volatility. We demonstrate an example of this complex dependency of vega on the volatility in Figure 2(b). Intuitively, when $\sigma>0$ is small, a tiny increase in volatility will lead to more trading and, therefore, fees collected by the LPs, i.e., a positive vega. On the other hand, when $\sigma>0$ is already very large, a further increase in volatility will, also, increase the probability of the event that the price collapses to zero; this leads to a drop in the price of the token $P$ which results in a negative vega.
@@
Figure 2: Valuation and vega with $P_{0}=1$, $\gamma=5bps$ (i.e., $\hat{\gamma}=5.0025bps$), $r=5\%$ (annualized) and $\Delta t=2$ seconds.